USE A PROPORTION TO SOLVE THE PROBLEM ABOUT DIRECT VARIATION.
A person's hair length varies directly with the number of weeks it has been growing. If Fred shaved his head bald and his hair was 0.15 inches long after 5 weeks, how long would it be afteran additional 4 weeks? Round to the nearest hundredth of an inch.

Answer by stanbon(75887)

A person's hair length varies directly with the number of weeks it has been growing. If Fred shaved his head bald and his hair was 0.15 inches long after 5 weeks, how long would it be after an additional 4 weeks? Round to the nearest hundredth of an inch.
---------------------------------
length = k*weeks
0.15 = k*5
k = 0.03
length = 0.03*weeks
--------------------------
length = 0.03*10
length = 0.3 inches
---------------------------
OR, using proportions:
x/10 = 0.15/5
x = 10(0.15/5)
x = 1.5/5 = 0.3 inches
========================
Cheers,
Stan H.